OpenStudy (anonymous):

An artist used silver wire to make a square that has a perimeter of 40 inches. She then used copper wire to make the largest circle that could fit in the square. How many more inches of silver wire did the artist use compared to the copper wire? (use 3.14 for pi)

OpenStudy (janu16):

Each side of the square has a length of \[40*\frac{ 1 }{ 4 } = 10\] inches The ardius of the circle is \[\frac{ 10 }{ 2 }\] = 5 inches so the circumference of the circle is 2* pi*5=10*3.14= 31.4 inches The perimeter of the square minus the circumference of the circle is 40 - 3.14 = 8.6 inches.

OpenStudy (janu16):

@Leylawashere

OpenStudy (anonymous):

Thank you so much!

OpenStudy (janu16):

np